Suppose that the production function of the firm is:

Q = 100L1/2.K1/2

 K= 100, P = $1, w = $50 and r = $40. Determine the quantity of labor that the firm should hire in order to maximize the profits. What is the maximum profit of this firm?


1
Expert's answer
2021-05-05T13:35:03-0400

Solution:

Determine the quantity of labor that the firm should hire in order to maximize the profits:


The production function of the firm is given by;

"Q=100L ^{0.5}K ^{0.5}"

Substitute K:

"Q=100L ^{0.5}(100) ^{0.5}"

"Q=100L ^{0.5}(10)"


"Q=1000L ^{0.5}"


R = P"\times" Q

R = "1\times 1000L ^{0.5} = 1000L ^{0.5}"


R = "1000L ^{0.5}"


Calculate the MPL:

This is the derivative of Labor to Revenue:

MPL = "\frac{\partial L} {\partial R} = \frac{500} {L ^{0.5}}"


Derive Total Cost function:

TC = wL + rK

TC = 50L + 40K

TC = 50L + 40(100)

TC = 50L + 4000

Find the derivative of the cost function with respect to Labor:

= 50

Now, set marginal product equal to zero and solve for L:

="\frac{500} {L ^{0.5}} -50 = 0"


="\frac{500} {L ^{0.5}} =50"


"50L ^{0.5} =500"

Square both sides:

2500L = 250000

L = 100


The quantity of labor that the firm should hire in order to maximize the profits = 100


Therefore:

Profit = TR - TC

Profit = "(1000L ^{0.5}) - (50L + 4000)"

Profit = "(1000(100) ^{0.5}) - (30(100) + 4000)"


Profit = "50,000 - 7,000 = 43,000"


The maximum profit of this firm = "\$43,000"